This is a PSA Question .. { Problem Solving Assesment }
 
Raman bought 5 flower pots of three different sizes for a total of Rs 290. Each pot cost 50 percentage more than the next one below its size . What is the largest flower pot ‘s cost ?
 
Pls detail explaination 

Grade:9

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Hi,

Assume cost of smallest size pot = x
So, cost of middle sized pot = 1.5x and their no. = a
and, cost of largest size pot = 2.25x and their no. = b

so,no. of smallest size pot = 5-a-b

Now, (5-a-b)x + a1.5x + b 2.25x = 290
x (5+0.5a+2.25b) = 290 = 29 * 10

So, x=29 and(5+0.5a+2.25b) = 10 ( because a and b are less than 5)

So, a=1, b=2
So cost of largest pot = 2.25 x = 65.25

thanks.